OpenStudy (anonymous):

What is the slope of the line that passes through (2, 12) and (4, 20)?

OpenStudy (anonymous):

\[\text{slope} = \frac{y_2 - y_1}{x_2 - x_1}\]

HanAkoSolo (jamierox4ev3r):

the answer is 4x

OpenStudy (anonymous):

from 2 to 4 is right 2 units from 12 to 20 is up 8 right 2, up 8,slope is \[\frac{8}{2}=4\]

HanAkoSolo (jamierox4ev3r):

in a y=mx+b, you get Y=4x

HanAkoSolo (jamierox4ev3r):

and you use abb50's formula

OpenStudy (anonymous):

I thought so just decided to post it just in case! thankyou all!!

OpenStudy (anonymous):

you can use a formula or you can think either way \[\frac{20-12}{4-2}=\frac{8}{2}=4\] no matter which way you do it